[figure: 9.7 cm]

Spin waves in La_{2}CuO_{4}

N. M. R. Peres and M. A. N. Araújo

Departamento de Física, Universidade de Évora, Rua Romão Ramalho, 59, P-7000-671

Évora, Portugal

and Centro de Física da Universidade do Minho, Campus Gualtar, P-4700-320 Braga, Portugal

We calculate the antiferromagnetic spin wave dispersion in a half-filled Hubbard model for a two-dimensional square lattice, and find it to be in excellent agreement with recent high-resolution inelastic neutron scattering performed on La_{2}CuO_{4} [Phys. Rev. Lett. 86, 5377 (2001)]. We also study the effect of different relation dispersions of the spin wave spectrum.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : N. M. R. Peres

Address for correspondence : Departamento de Física, Universidade de Évora, Rua Romão Ramalho, 59, P-7000-671 Évora, Portugal

Email address : peres@uevora.pt